OpenStudy (anonymous):

Please help me! Will give medal :) Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of 14 inches and a height of 4 inches. Each leg is 4 inches by 4 inches by eighteen inches. How much area will Evan have to paint? Use three point one four for pi. Show your work.

OpenStudy (anonymous):

@Ahlpa_Omega32

OpenStudy (anonymous):

please help i have this and 2 more questions to finish my class!!

OpenStudy (bibby):

The area of the stool is equal to the area of the cylinder seat + 4 leg areas. First find the length area. You have a rectangular prism measuring 4x4x18 what's the area of that figure?

OpenStudy (anonymous):

288

OpenStudy (bibby):

So we have the area of the legs. Now what is the area of a cylinder with diameter of 14 inches and a height of 4 inches Remember that the area of a cylinder is pi * r^2 * h

OpenStudy (bibby):

that might not be it actually

OpenStudy (anonymous):

umm.. what? ... the area is 615.44

OpenStudy (bibby):

yeah that's wrong. find the surface area

OpenStudy (bibby):

how'd you get that number?

OpenStudy (anonymous):

3.14*radius^2*h... like u said

OpenStudy (bibby):

I said those weren't the right formulas. If we're painting the chair, we have to use surface area formulas

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

Find the surface area formulas for a rectangular prism and cylinder and apply them

OpenStudy (anonymous):

rectangular prism:2(wl+hl+hw)... cylinder:2πrh+2πr2

OpenStudy (bibby):

w-4 l-4 h18

OpenStudy (anonymous):

ok hold on :)

OpenStudy (anonymous):

for rectangular prism its:872........

OpenStudy (bibby):

I got 320 2(16+72+72)

OpenStudy (bibby):

why would I do your homework? I'm here to give you a guiding hand not give you the answers dear

OpenStudy (anonymous):

yea ... but idk i think my calculator is damaged lol.... hold on, so you got 320 for rectangular prism.. so let me do the cylinder

OpenStudy (bibby):

2πrh+2πr2 remember that pi is approximated to 3.14 r = 14/2 = 7 h = 4

OpenStudy (anonymous):

there u go.... i thought the height was 18.. lol

OpenStudy (bibby):

that was for the rectangle

OpenStudy (anonymous):

ehh.... i need to learn.... hold on

OpenStudy (anonymous):

cylinder: 7817.8464

OpenStudy (bibby):

lemme see \[\huge 2πrh+2πr^2\]\[\huge 2π7*4+2π7^2\]\[\huge 28*2π+2π49\]\[\huge 56π+98π\]154*3.14=483.56

OpenStudy (anonymous):

oh ok... so do i add up the to answers?

OpenStudy (bibby):

not exactly. Logically this is what the chair looks like: